Hathaway Health Club sold three-year memberships at a reduced rate during its opening promotion. It sold 1,000 three-year nonrefundable memberships for $357 each. The club expects to sell 100 additional three-year memberships for $900 each over each of the next two years. Membership fees are paid when clients sign up. The club's bookkeeper has prepared the following income statement for the first year of business and projected income statements for Years 2 and 3.

Cash-basis income statement:
Year 1 Year 2 Year 3
Sales $357,000 $90,000 $90,000
Equipment* $106,000 $0 $0
Salaries and wages 49,000 49,000 49,000
Advertising 4,810 4,810 4,810
Rent and utilities 35,020 35,020 35,020
Net income (loss) $162,170 $1,170 $1,170
*Equipment was purchased at the beginning of year 1 for $106,000 and is expected to last for three years and then to be worth $1,060.
Required:
Convert the income statements for each of the three years to the accrual basis. Indicate a net loss with a minus sign.
Hathaway Health Club
Income Statements
Year 1 Year 2 Year 3
Sales $ $ $
Expenses:
Depreciation $ $ $
Salaries and wages
Advertising
Rent and utilities
Total expenses $ $ $
Net income (loss) $ $ $
Answer #1

Answer:

Year 1 Year 2 Year 3

Sales

119,000

149,000

179,000

Expenses:

Depreciation

34,980

34,980

34,980

Salaries and wages

49,000

49,000

49,000

Advertising

4,810

4,810

4,810

Rent and utilities

35,020

35,020

35,020

Total expenses

123,810

123,810

123,810

Net Income (loss)

(4,810)

25,190

55,190

Calculation:

Here, we need to find the net Income (loss) for three years in the accrual basis. So first we need to divide the total sales with the number of years to calculate the sales in the first year.

Sales in the 1st Year = Total Sales/Number of Years = 357,000 / 3 = 119,000

Then we need to divide the total sales by the number of years, and add it to the first year's sales to calculate the sales in the second year.

Sales in the 2nd Year = Total Sales /Number of Years+ Sales in the 1st Year = 90000 / 3 +119,000 = 149,000

Then we need to divide the total sales by the number of years, and add it to the second year's sales to calculate the sales in the third year.

Sales in the 2nd Year = Total Sales /Number of Years+ Sales in the 2nd Year = 90000 / 3 + 149,000 = 179,000

Then we need to find the depreciation value. For that we need to subtract salvage value of assets from cost of assets, then divided by the useful life

Depreciation = Useful Life Cost of Assets − Salvage Value = (106,000 − $1,060)/3 = $34,980

Rest of the expenses remains the same. So we need to total the expenses for the three years respectively and deduct from the sales to get the net income/loss.